# On local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms

### v2

@@ 3. Rigidity theorem for local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms
Let $(X,g)$ be a complex $n$-dimensional Hermitian manifold with the Hermitian form $\omega$, namely, $\omega=g(J\cdot,\cdot)$ with $J$ being the induced almost complex structure on $(X,g)$. For simplicity, in the present article we write $(X,\omega)$ to indicate that $\omega$ is the Hermitian form. In terms of local holomorphic coordinates $(z_{1},\ldots,z_{n})$ on $X$, we may write $\omega=\sqrt{-1}\sum_{i,j=1}^{n}g_{i\overline{j}}dz_{i}\wedge d\overline{z_{j}}$ locally. Then, for $1\leq p\leq n$ we have
@@ 4.1. The unit sphere bundles on Hermitian manifolds
Now, we consider Case (b), i.e., $(N,\omega_{N})$ is a period domain equipped with the canonical invariant metric and $F$ is a horizontal holomorphic map, i.e., $F:U\to N$ is a holomorphic map such that $dF_{x}(T_{x}(M))$ lies inside the horizontal tangent space $T^{h}_{F(x)}(N)\subset T_{F(x)}(N)$ for all $x\in U$. Then, $F(U)\subset N$ is an integral manifold for the horizontal distribution on the period domain $N$. By a result of Carlson-Toledo [CT88, Proposition 5.3], we have $dF_{x}(T_{x}(M)),dF_{x}(T_{x}(M))]=0$ for $x\in U$. It follows from Peters [Pet90, Corollary 1.8] that the holomorphic bisectional curvature $R_{v\bar{v}\xi\bar{\xi}}(N,\omega_{N})\leq 0$ along the directions $v,\xi\in dF_{x}(T_{x}(M))$, $x\in U$. Hence, we still have $(\partial\overline{\partial}\rho_{2})(f(w_{0},v_{0}))(f_{*}(\eta),\overline{f_{*}(\eta)})$ $\geq$ 0. By the same argument as in the above, this contradicts with (4.2), and the proof of the second assertion is complete. ∎
@@ 4.2. The Umehara algebra on bounded symmetric domains and its applications
$\begin{split}\widetilde{\phi}_{m}(\zeta)=&\det\begin{bmatrix}{\lambda_{1}\over(1-|\zeta|^{2})^{2}}&&&&&\\ &{\lambda_{1}\over 1-|\zeta|^{2}}{\bf I}_{p_{1}}&&&&\\ &&\lambda_{1}{\bf I}_{m_{1}-p_{1}-1}&&&\\ &&&\lambda_{2}{\bf I}_{m_{2}}&&\\ &&&&\ddots&\\ &&&&&\lambda_{k}{\bf I}_{m_{k}}\end{bmatrix}\\ =&{\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{p_{1}+2}},\end{split}$

### v3

@@ 3. Rigidity theorem for local holomorphic maps between Hermitian manifolds preserving $(p,p)$-forms
Let $(X,g)$ be a complex $n$-dimensional Hermitian manifold with the Hermitian form $\omega$, namely, $\omega=g(J\cdot,\cdot)$ with $J$ being the induced almost complex structure on $(X,g)$. For simplicity, in the present article we write $(X,\omega)$ to indicate that $\omega=:\omega_{X}$ is the Hermitian form. In terms of local holomorphic coordinates $(z_{1},\ldots,z_{n})$ on $X$, we may write $\omega=\sqrt{-1}\sum_{i,j=1}^{n}g_{i\overline{j}}dz_{i}\wedge d\overline{z_{j}}$ locally. Then, for $1\leq p\leq n$ we have
@@ 4.1. The unit sphere bundles on Hermitian manifolds
Now, we consider Case (b), i.e., $(N,\omega_{N})$ is a period domain equipped with the canonical invariant metric and $F$ is a horizontal holomorphic map, i.e., $F:U\to N$ is a holomorphic map such that $dF_{x}(T_{x}(M))$ lies inside the horizontal tangent space $T^{h}_{F(x)}(N)\subset T_{F(x)}(N)$ for all $x\in U$. As in Case (a), we may assume without loss of generality that $p=m$ so that $F^{*}\omega_{N}^{m}=\lambda\omega_{M}^{m}$ holds, and thus $F$ is a holomorphic immersion. Then, $F(U)\subset N$ is an integral submanifold for the horizontal distribution on the period domain $N$. By a result of Carlson-Toledo [CT88, Proposition 5.3], we have $dF_{x}(T_{x}(M)),dF_{x}(T_{x}(M))]=0$ for $x\in U$. It follows from Peters [Pet90, Corollary 1.8] that the holomorphic bisectional curvature $R_{v\bar{v}\xi\bar{\xi}}(N,\omega_{N})\leq 0$ along the directions $v,\xi\in dF_{x}(T_{x}(M))$, $x\in U$. Hence, we still have $(\partial\overline{\partial}\rho_{2})(f(w_{0},v_{0}))(f_{*}(\eta),\overline{f_{*}(\eta)})$ $\geq$ 0. By the same argument as in the above, this contradicts with (4.2), and the proof of the second assertion is complete. ∎
@@ 4.2. The Umehara algebra on bounded symmetric domains and its applications
$\tiny\widetilde{\phi}_{m}(\zeta)=\det\begin{bmatrix}{\lambda_{1}\over(1-|\zeta|^{2})^{2}}&&&&&\\ &{\lambda_{1}\over 1-|\zeta|^{2}}{\bf I}_{p_{1}}&&&&\\ &&\lambda_{1}{\bf I}_{m_{1}-p_{1}-1}&&&\\ &&&\lambda_{2}{\bf I}_{m_{2}}&&\\ &&&&\ddots&\\ &&&&&\lambda_{k}{\bf I}_{m_{k}}\end{bmatrix}={\lambda_{1}^{m_{1}}\cdots\lambda_{k}^{m_{k}}\over(1-|\zeta|^{2})^{p_{1}+2}},$